From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Feb 27 2003 - 15:50:43 EST
------------------------------------------------------------ Date: Thu Feb 27 15:50:42 2003, Shift: 08:00-16:00 Supervisor: bhs, with: Log book#14 pages 38 - 38 Shift Summary: ------------------------------------------------------------ Shift report B. H. Samset This morning when I came on shift, the sun was shining bright. A Cooling Failure had just been fixed that had troubled the machine all night. At ten'o'clock we were promised beam, though they did not say just when, but at eleven they sadly reported a Cooling Failure once again... Not everything was gloomy though, there was still reason to smile, 'cause this problem meant that our IR was open for a while. So Dana and Hiro saddled up and went in to do the chore of cabling up the phototubes of our new pride: C4. Now, because I'm such a good shifter, and because it's so much fun, I sat down at the DAQ to do a TPC pedestal run. This turned out to be somewhat hard, no data at all got through, and Flemming had to come up and recycle the power on TPM2. I then had a look at the pedestals in T1s first two padrows, but as I click the buttons, by gosh, no time bin profile shows! It appeares these rows were both quite dead, and we found out how long for: It happened sometime in sequence 5 of run 8044. Luckily another power cycle got the sleeping rows to awaken, now everything seems a-ok, and new pedestals have been taken. At three: They're promising imminent beam, current problems are dispersed, though they will run a 6 bunch ramp to test Landau cavities first. That's it for now, and if you find this report somewhat shrouded in fog, I advice you to go on the net and check out our nice e-log. A note of caution at the end: when I came in the coff- ee maker had been on all night - please remember to turn it off! ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Thu Feb 27 2003 - 15:56:32 EST